# LoopServe: An Adaptive Dual-phase LLM Inference Acceleration System for Multi-Turn Dialogues

### v1

@@ 2.2. Efficient Long-Context Inference
2.2.1. Context Compression Methods
Context compression strategies aim to reduce effective sequence length by systematically eliminating non-essential content or distilling key information from verbose inputs. These approaches effectively transform lengthy inputs into more computationally manageable representations. Firstly, the filtering-based methods, such as LLMLingua (Jiang et al., 2023), LLMLingua-v2 (Pan et al., 2024), and CompAct (Yoon et al., 2024), selectively identify and preserve high-relevance content, enabling models to process critical information without engaging the entire sequence at maximum resolution. Alternatively, RAG-based techniques (Zhao et al., 2024; Jiang et al., 2024b; Wang et al., 2025; Chen et al., 2025; Edge et al., 2024) employ a two-stage process: initially constructing knowledge graphs through extraction of essential semantic triples from input text, followed by synthesizing this structured knowledge into condensed representations that serve as reduced-dimension inputs for LLM processing. While these compression strategies substantially reduce computational and memory requirements during inference, they inevitably sacrifice some degree of granular detail, potentially compromising the fidelity and precision of LLM-generated outputs.
2.2.2. KV-based Optimization
To mitigate the computational burden, KV-based approaches focus on reducing the number of attention weight calculations during both prefilling and decoding stages, as summarized in Table 2.
Prefilling-stage Optimization. Given a input $X_{i}$ with $n_{i}$ length, the prefilling stage requires $O(n^{2})$ computations to compute each attention matrix across all token pairs, which is a significant computational bottleneck. Recent approaches such as Minference (Jiang et al., 2024a), FlexPrefill (Lai et al., 2025), and CritiPrefill (Lv et al., 2024) address this by selectively discarding negligible attention values, effectively sparsifying each attention head matrix:
$\min\,\left|\mathbf{{A}}^{j}_{i}-\mathbf{\hat{A}}^{j}_{i}\right|,s.t.\ ,\mathbf{\hat{A}}^{j}_{i}=\textsf{Softmax}\left(\frac{\mathbf{Q}^{j}(\mathbf{K}^{j})^{\top}}{\sqrt{d_{k}}}-c(1-\mathbf{M})\right),$
where $\mathbf{M}\in\{0,1\}^{n\times n}$ represents a binary sparsity mask, where $\mathbf{M}[a][b]=1$ indicates that the attention weight $\mathbf{A}^{j}_{i}[a][b]$ should be preserved. The constant $c=10^{8}$ ensures that masked positions ($\mathbf{M}[a][b]=0$) receive negligible attention weight after softmax normalization. This optimization reduces the computational complexity from $O(n^{2})$ to $O(\alpha n^{2})$ per head, where the sparsity factor $\alpha\ll 1$. Specifically, Minference (Jiang et al., 2024a) improves efficiency by selecting key vertical and slash attention patterns or important blocks, while FlexPrefill (Lai et al., 2025) enhances this approach by dynamically adjusting the sparsity rate based on attention recovery rates. In contrast, CritiPrefill (Lv et al., 2024) partitions keys and queries into blocks and simply selects the top-$k$ most important key blocks for computation. However, as shown by our motivational experiments (Section 3), the sparsity and importance patterns in attention matrices are highly dynamic and input-dependent, varying significantly across turns, heads, and layers. Consequently, fixed-pattern selection approaches often fail to achieve robust performance in real multi-turn scenarios.
Decoding-stage Optimization. During autoregressive text generation, LLMs utilize KV caches to store and reuse partial previously computed Key and Value embeddings, thereby reducing computational redundancy. Firstly, token-level selection approaches (Ge et al., 2024; Li et al., 2024b), such as , H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), aim to select important tokens from inputs before output generation. These methods assume that critical input tokens required for generating outputs can be identified beforehand and that these tokens significantly overlap with those in the last observation window of the inputs. As analyzed in Section 3, these approaches perform well on existing benchmarks like LongBench (Bai et al., 2024a, b), because such benchmarks place user queries at the end of the input sequence. This assumption holds in existing benchmarks, where user queries are always placed at the end of the input sequence. In such settings, token-level selection methods perform well: by focusing on the most recent tokens (which include the user query), these methods can accurately preserve the necessary context for high-quality output, while significantly reducing memory and computation.
However, this benchmark design does not match real-world multi-turn dialogues, where user queries can appear anywhere in the input. When queries are not always at the end, token-level selection methods like SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) often miss the critical tokens needed for accurate output, resulting in lower performance and poor generalization. As shown in Section 3.2, their effectiveness drops significantly when queries are in the middle or at the front of the context. This highlights the need for adaptive, context-aware KV selection methods that can handle dynamic query positions and dependencies in practical dialogue scenarios.
2.3. Long-Context Benchmarks
Recently, various long-context benchmarks (Li et al., 2024c, 2025; Hsieh et al., 2024; Kim et al., 2025; An et al., 2023; Qin et al., 2024) have been proposed to evaluate LLMs, such as NumericBench (Li et al., 2025), LongBench (Bai et al., 2024a, b), and LongEval (Dacheng Li* and Zhang, 2023). However, these benchmarks have two significant limitations that hinder their applicability to real-world scenarios. First, they assume that user queries are always positioned at the end of the input context. This architectural bias simplifies the evaluation process but fails to capture the complexity of real-world interactions, where queries may appear at the beginning, middle, or interspersed throughout the input. As a result, existing KV-compression approaches perform well on these benchmarks because they are optimized for scenarios where the user question is placed at the end of the input. However, these methods struggle to generalize to real-world human-LLM interactions. Second, most of these benchmarks are single-turn, focusing on isolated question-answering or summarization tasks. This approach overlooks the multi-turn nature of real-world conversations, where contextual dependencies across multiple turns are critical (LI et al., 2025).
To address these limitations, we propose a multi-turn benchmark with 11 datasets in Section 4 with diverse query positions and interaction patterns, providing a more realistic evaluation of LLMs for long-context tasks.
@@ 2.3. Long-Context Benchmarks
Table 2. LLM acceleration model comparisons, following (LI et al., 2025). P and D denote whether the model has optimization in the Prefilling and Decoding phases, respectively. $n$ is the token size of the input, $m$ is the generation token size, and $c$ and $k$ are constants with $c,k\ll n$ and $c,k\ll m$.
@@
LLMLingua (Pan et al., 2024) $\times$ $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
A-shape (Xiao et al., 2024) $\times$ $O(n)$ $O(kn)$ $O(nm)$
Tri-shape (LI et al., 2025) $\times$ $O(n)$ $O(kn)$ $O(nm)$
MInference (Jiang et al., 2024a) $\times$ $O(n)$ $O(kn)$ $O(nm)$
SLLM (Xiao et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
SnapKV (Li et al., 2024b) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
AdaKV (Feng et al., 2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
LoopServe $O(k)$ $O(kn)$ $O(k(m-c)+nc)$
3. Motivational Experiments and Insights
To better understand the fundamental challenges of accelerating LLM inference in multi-turn dialogues, we conduct a series of motivational experiments. These experiments, presented in Sections 3.1 and 3.2, provide key observations on the dynamic sparsity of attention patterns and the impact of query position on acceleration effectiveness. Based on these insights, we introduce the LoopServe system in Section 5, which is specifically designed to address these unique challenges in real-world multi-turn dialogue scenarios.
3.1. Key Point 1: Uncertain Attention Patterns
As discussed in Section 2, each attention head matrix is highly sparse. Current approaches (Jiang et al., 2024a; Xiao et al., 2024; Feng et al., 2024; LI et al., 2025; Li et al., 2024b) for accelerating the prefilling and decoding phases often rely on attention matrix sparsification and important KV token selection. However, these methods are built on a strong assumption: the attention sparsification in during the prefilling and important tokens in decoding follow fixed patterns, and their positions can be identified offline. However, the sparsity patterns of attention matrices are dynamic and unpredictable, varying across different inputs, heads, and layers. This variability challenges the effectiveness of fixed-pattern approaches. In this subsection, we first reveal the sparsity of each attention head and then demonstrate that the sparsity patterns vary across different attention heads and input sequences.
(a) All heads of Lamma 3.
(b) Average ratio among all heads.
Figure 1. Attention head sparsity. Only a small fraction of vertical and slash lines in each attention head of LLMs can recover most of the attention weights, revealing strong sparsity in the attention matrix.
3.1.1. Motivational Observation 1: only 90% of vertical lines and slash lines can collectively account for most (e.g., 90%) of the attention weight.
In the attention matrix, a vertical line is a single column, showing that one token is attended to by all other tokens in the sequence. This often happens with special tokens like separators or keywords. A slash line is a diagonal, showing that each token mainly focuses on its nearby tokens, which is a pattern of local attention. We conduct experiments on a SAMSum QA (Bai et al., 2024a) dataset $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{|\mathcal{D}|}$ using the Llama-3.1-8B-Instruct model with $n_{h}$ attention heads. For each user query $X_{i}$ of length $n_{i}$, each $k$-th attention matrix $\mathbf{A}^{k}_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ contains $n_{i}$ slash lines $\mathcal{S}^{k}_{i}$ and $n_{i}$ vertical lines $\mathcal{V}^{k}_{i}$. The total attention weight for each slash line $s^{k}\in\mathcal{S}^{k}_{i}$ is computed as $\sum_{(a,b)\in s^{k}}\mathbf{A}^{k}_{i}[a][b]$, and the total attention weight for each vertical line $v^{k}\in\mathcal{V}^{k}_{i}$ is $\sum_{(a,b)\in v^{k}}\mathbf{A}^{k}_{i}[a][b]$. Next, we only use the top $\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{S}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$. Therefore, for each $k$-the attention head, the ratio of the attention weight within these selected lines to the total attention weight can be computed as follows: $r^{k}_{i}=\frac{1}{|\mathcal{D}|}\sum_{X_{i}\in\mathcal{D}}\frac{\sum_{(a,b)\in\mathcal{\hat{S}}^{k}_{i}\cup\mathcal{\hat{V}}^{k}_{i}}\mathbf{A}^{k}_{i}[a][b]}{n_{i}}.$ Also, the average ratio among $n_{h}$ attention heads can be computed as $\frac{1}{n_{h}}\sum_{j=1}^{n_{h}}r^{k}_{i}$. As shown in Figure 1 (a), increasing $\eta$ raises the accumulated attention weight across different heads; a small percentage of slash and vertical lines captures most of the attention weight. Figure 1 (b) shows that both LLaMA and Mistral exhibit this sparsity: covering 90% of the total attention weight requires selecting only 10% of slash and vertical lines. This indicates that, due to the sparsity in attention heads, vertical and slash lines are fundamental for efficient selection.
3.1.2. Motivational Observation 2: The positions of the top vertical and slash lines within the same head vary across different user inputs.
Given the LLM model $M_{\theta}$ with $n_{h}$ attention heads, the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, and the percentage $\eta$, we select the top-$\eta\cdot 2n_{i}$ slash lines $\mathcal{\hat{S}}^{k}_{i}\subseteq\mathcal{{S}}^{k}_{i}$ and vertical lines $\mathcal{\hat{V}}^{k}_{i}\subseteq\mathcal{V}^{k}_{i}$ for each user query $X_{i}$ under each $k$-th head. For two queries $X_{i}$ and $X_{j}$, we compute the line overlap of the selected vertical and slash lines as: $r^{k}_{i,j}=\frac{|\mathcal{\hat{S}}_{i}^{k}\cap\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cap\mathcal{\hat{V}}_{j}^{k}|}{|\mathcal{\hat{S}}_{i}^{k}\cup\mathcal{\hat{S}}_{j}^{k}|+|\mathcal{\hat{V}}_{i}^{k}\cup\mathcal{\hat{V}}_{j}^{k}|}.$ The average important line overlap ratio across $n_{h}$ heads can then be computed as $\frac{1}{n_{h}|\mathcal{D}|^{2}}\sum_{k=1}^{n_{h}}\sum_{X_{i},X_{j}\in\mathcal{D}}r^{k}_{i,j}$. In particular, we use SAMSum QA (Bai et al., 2024a) dataset along with Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3, and vary $\eta\in\{0.1,0.15,0.2,0.25,0.3\}$. As shown in Figure 2, the overlap of important lines across both LLMs is mostly below 0.5 for different inputs in most heads. This shows that different texts produce distinct line patterns, even within the same attention head. Therefore, it is not practical to identify important lines offline and use them directly for online inference.
(a) Llama-3.1-8B-Instruct.
(b) Mistral-7B-Instruct-v0.3.
Figure 2. Line overlap of the same head with different inputs. The important attention lines differ significantly across inputs, even for the same attention head, highlighting the dynamic and input-dependent nature of attention sparsity.
(a) Attention map.
(b) Important line overlap.
Figure 3. Different patterns in the same head. It illustrates that important attention lines differ even within different segments of the same input, emphasizing the locality and unpredictability of attention patterns.
3.1.3. Motivational Observation 3: Given an input $X_{i}=[C_{i}^{1},C^{2}_{i}]$, where $X_{i}$ consists of two sequential segments $C_{i}^{1}$ and $C_{i}^{2}$. The important positions of the top vertical and slash lines within the same head differ between $C^{1}_{i}$ and $C^{2}_{i}$. Both $C^{1}_{i}$ and $C^{2}_{i}$ exhibit their own local attention sparsity patterns.
As shown in Figure 3 (a), the prominent slash and vertical lines in the attention matrix of $C^{1}$ do not appear in $C^{2}$. Instead, the attention part of ${C^{2}}$ exhibits distinct local patterns. To qualitatively verify this, we use SAMSum QA (Bai et al., 2024a) data with Llama-3.1-8B-Instruct to conduct experiments. Given the dataset $\mathcal{D}=\{X_{i}\}_{i=1}^{|\mathcal{D}|}$, each $X_{i}$ is split into two sequential parts as $X_{i}=[C^{1}_{i},C^{2}_{i}]$. We denote the attention matrix of $C^{1}_{i}$ as $\mathbf{A}^{k}_{C^{1}_{i}}\in\mathbb{R}^{|C^{1}_{i}|\times|C^{1}_{i}|}$ and the attention matrix of $C^{2}_{i}$ as $\mathbf{A}^{k}_{C^{2}_{i}}\in\mathbb{R}^{|C^{2}_{i}|\times n_{i}}$. Next, we select the top-$\eta$ slash and vertical lines for $\mathbf{A}^{k}_{C^{1}_{i}}$ and $\mathbf{A}^{k}_{C^{2}_{i}}$, denoted as $L^{k}_{C^{1}_{i}}$ and $L^{k}_{C^{2}_{i}}$, respectively. The overlap rate of importat lines between $C^{1}_{i}$ and $C^{2}_{i}$ as $r^{k}_{C_{i}^{1}\to C_{i}^{2}}={\sum_{l\in L^{k}_{C^{1}}}\mathbb{I}(l\in L^{k}_{C^{2}})}/{|L^{k}_{C^{2}}|},$ where $\mathbb{I}(l\in L^{k}_{C^{2}})=1$ if the line $l\in L^{k}_{C^{2}}$, indicating that an important line $l$ of $C_{i}^{1}$ is still important in $C_{i}^{2}$. Then, given data $\mathcal{D}$ and the $k$-th attention head, the average overlapping ratio is $\frac{1}{\mathcal{D}}\sum_{X_{i}\in\mathcal{D}}{r^{k}_{C_{i}^{1}\to C_{i}^{2}}}$. As shown in Figure 3 (b), under different $\eta$, The overlap rate between contexts in a user query remains low and highly volatile, indicating distinct local attention patterns for different subsections of input $X_{i}$.
This observation directly highlights that it is challenging to use a subsection of the input to determine the patterns of the entire input, as distinct local patterns exist within different subsections. Consequently, existing prefilling acceleration approaches, such as Minference (Jiang et al., 2024a), which rely on the last window (e.g., $C^{2}_{i}$ )to determine the patterns for the whole input $X_{i}$ and select important attention blocks, fail to achieve consistent performance in real-world applications. Similarly, decoding acceleration approaches, such as SnapKV (Li et al., 2024b), H2O (Zhang et al., 2023b), and Keyformer (Adnan et al., 2024), which use the last observation window (e.g., last several tokens of $C_{i}^{1}$) to identify important tokens for generating outputs (e.g., the second context $C_{i}^{2}$), are also unreliable. As a result, these methods struggle to deliver good performance in practical scenarios.
3.2. Key Point 2: Question Position Matters.
Based on the above motivational experiments, both prefilling-based and decoding-phase acceleration approaches, that rely on offline sparse pattern discovery or fixed sparse patterns, should incapable of performing well in real-world use cases. However, the results reported in their papers are often comparable to original LLMs using the full attention matrix without any compression. Why? The reason is that existing long-context benchmarks, such as Longbench (Bai et al., 2024a, b), place each user question $q_{i}$ at the end of the input sequence $X_{i}=[C_{i},q_{i}]$. Then, the LLMs will answer the question $q_{i}$ based on the context $C_{i}$. In this setup, by relying only on the last observation window of the input (i.e., the user questions), existing acceleration approaches identify important tokens in the context that are highly correlated with the user’s specific questions. This partially alleviates the uncertainty caused by distinct local patterns observed in real-world scenarios.
3.2.1. Motivational Observation 4: We reveal that relying solely on the last observation window of the input cannot reliably identify the important tokens from the input $X_{i}$ for generating the output $Y_{i}$.
Existing approaches, such as H2O (Zhang et al., 2023b), SnapKV (Li et al., 2024b), and AdaKV (Feng et al., 2024), propose selecting the top-$B$ most important tokens $\hat{X}_{i}\subseteq X_{i}$ from the input $X_{i}$ before generating the output tokens. Given the input $X_{i}$, we denote the last observation window of $X_{i}$ with length $n_{s}$ as $X^{obs}_{i}=X_{i}[n_{i}-n_{s}:n_{i}]$. Under the LLM model $M_{\theta}$ with $n_{h}$ attention heads, these approaches select the top-$B$ tokens $\hat{X}_{i}\subseteq X_{i}$ where $|\hat{X}_{i}|=B$ as follows:
(4) $\hat{X}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in X^{obs}_{i}}\mathbf{A}^{k}_{i}[a][b],$
Then, given the output $Y_{i}=[x_{n_{i}+1},x_{n_{i}+2},\cdots,x_{n_{i}+n_{o}}]$ corresponding to the input $X_{i}$, the ground-truth top-$B$ important tokens $X^{*}_{i}\subseteq X_{i}$ can be obtained as follows:
(5) $\hat{X}^{*}_{i}=\arg\max_{\hat{X}_{i}\subseteq X_{i}}\sum_{j=1}^{n_{h}}\sum_{a\in\hat{X}_{i}}\sum_{b\in Y_{i}}\mathbf{A}^{k}_{i}[a][b],$
The overlap between $\hat{X}_{i}$ and $\hat{X}^{*}_{i}$ can be computed as $r_{i}=\frac{|\hat{X}_{i}\cap\hat{X}^{*}_{i}|}{B}$ and the average rate regarding the data $\mathcal{D}$ is $\frac{\sum_{X_{i}\in\mathcal{D}}{r_{i}}}{|\mathcal{D}|}$. We set the budget $B=0.1|X_{i}|$, corresponding to 10% of the tokens in each input $X_{i}$. We use Llama-3.1-8B-Instruct and construct the dataset from LongEval’s topic retrieval set (Dacheng Li* and Zhang, 2023), where each instance contains three topic contexts and an instruction (framed as a question) to summarize two of them. For each instance $X_{i}=(C_{i},q_{i})$, we create three variants by placing the instruction or question $q_{i}$ at the beginning, middle, or end of the context $C_{i}$. As shown in Figure 4 (a), the average overlap of important tokens is highest when the question is at the end, but much lower when placed earlier, indicating that relying only on the last observation window misses relevant information unless the question appears last.
(a) Important token overlapping.
(b) Question in different position.
Figure 4. Question position matters. When user queries are not at the end of the input, both the identification of important input tokens and the task performance of KV compression methods decline, indicating their reliance on queries being at the end.
Also, as shown in Figure 4 (b), SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) perform similarly to the original LLM when the question is at the end of the input. However, their performance drops significantly when the question appears in the middle or at the beginning. It is because SnapKV (Li et al., 2024b) and AdaKV (Feng et al., 2024) rely on the last few tokens to select the most important tokens from the input. When the user questions are at the end of inputs, the last few tokens include the question itself, which helps identify tokens related to the question and generate high-quality outputs. In contrast, when the questions are placed in the beginning or middle, the last few tokens are unrelated to the questions and answers, making it difficult to identify the important tokens and leading to performance drop . This shows that current approaches depend heavily on the benchmark’s question placement, which does not reflect real-world input patterns. Consequently, they struggle to generalize to more realistic scenarios.
4. Multi-turn Long-Context Benchmarks
As discussed in Section 2.3, existing benchmarks mainly focus on single-turn tasks and place user queries at the end of the context, failing to reflect the complexity of real-world multi-turn dialogues. To address this, we propose a new multi-turn long-context benchmark that enables more comprehensive and realistic evaluation of LLMs in conversational scenarios, which is released in Huggingface.
Table 3. Multi-turn dataset statistics.
Type Dataset $|D|$Turn Avg Token Metric

### v2

@@ A.1 Important Notations Table
Table 2 provides detailed definitions of important notations appearing in this paper.
A.2 Summary Table of KV-based approaches
@@
Table 3: LLM acceleration model comparisons, following LI et al. (2025). P and D denote whether the model has optimization in the Prefilling and Decoding phases, respectively. $n$ is the token size of the input, $m$ is the generation token size, and $c$ and $k$ are constants with $c,k\ll n$ and $c,k\ll m$.
@@
LLMLingua Pan et al. (2024) $\times$ $O(\alpha n)$ $O(\alpha^{2}n^{2})$ $O(\alpha nm)$
A-shape Xiao et al. (2024) $\times$ $O(n)$ $O(kn)$ $O(nm)$
Tri-shape LI et al. (2025) $\times$ $O(n)$ $O(kn)$ $O(nm)$
MInference Jiang et al. (2024a) $\times$ $O(n)$ $O(kn)$ $O(nm)$
SLLM Xiao et al. (2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
SnapKV Li et al. (2024b) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
AdaKV Feng et al. (2024) $\times$ $O(k)$ $O(n^{2})$ $O(km)$
LoopServe $O(k)$ $O(kn)$ O(k(mc)+nc)O(k(m-c)+nc)
Table 3 summarizes the time complexity for each KV-based approach.
A.3 Supplementary Figure for Motivational Observation 2
Figure 6: Overlap rate of each head regarding different inputs of Mistral-7B-Instruct-v0.3
Figure 7: Ablation Study of LoopServe on Qwen2.5-7B-Instruct.
Figure 7 shows that for most heads, the overlap of Mistral-7B-Instruct-v0.3 remains below 0.5. Please refer to the detailed analysis in Section 3.1 motivational experiment 2.
A.4 Supplementary Results for Ablation Study
Figure 7 shows that LoopServe applied on Qwen2.5-7B-Instruct achieves the best performance, indicating the significance of both components. Please refer to the detailed analysis in ablation study in Section 6.2.
A.5 Long-context Multi-turn LongBench
Recently, various long-context benchmarks, such as NumericBench Li et al. (2025), LongBench Bai et al. (2024a), and LongEval Dacheng Li* & Zhang (2023) have been proposed to evaluate LLMs. However, these benchmarks have two main limitations: (1) They assume user queries always appear at the end of the input, which does not reflect real-world scenarios with queries at arbitrary positions. This bias favors KV-compression methods optimized for end-positioned queries, limiting their generalizability. (2) Most benchmarks are single-turn, overlooking the multi-turn dependencies crucial for realistic conversations LI et al. (2025).
To overcome these issues, we propose a multi-turn benchmark spanning 11 datasets with diverse query positions and interaction patterns, enabling more realistic long-context LLM evaluation.
A.5.1 The Design of Multi-turn LongBench
We represent each $m$-turn long-context data instance in our dataset using a structured format. Specifically, each data instance $\mathcal{I}_{i}$ consists of $m$ turns, where each turn contains a triplet of context, question, and answer. The complete dataset can be formally denoted as:
$\mathcal{D}=\{\mathcal{I}_{i}=[(C_{i,1},q_{i,1},a_{i,1}),(C_{i,2},q_{i,2},a_{i,2}),\dots,(C_{i,m},q_{i,m},a_{i,m})]\}_{i=1}^{|\mathcal{D}|},$
where $C_{i,j}$ is the context at the $j$-th turn of the instance $\mathcal{I}_{i}$, which can be empty, $q_{i,j}$ is the corresponding user question, and $a_{i,j}$ denotes the generated answer of the LLMs for $q_{i,j}$. We design diverse formats for each multi-turn long-context data instance as follows:
A.5.2 Multi-turn LongBench Generation
Table 4: Multi-turn dataset statistics.
Type Dataset $|D|$ #Turn Avg Token Metric